A boy pulls Hus toy on a smooth horizontal surface with a rope inclined at 60° to the horizontal. If the effective force pulling the toy along the horizontal surface is 5N, calculate the tension in the rope


1
Expert's answer
2021-03-08T08:10:44-0500

We can find the tension in the rope from the geometry:


"cos\theta=\dfrac{T_x}{T},""T=\dfrac{T_x}{cos\theta}=\dfrac{5\ N}{cos60^{\circ}}=10\ N."